# Hierarchical Long Video Understanding with Audiovisual Entity Cohesion and Agentic Search

### v1

@@ 1 Introduction
The rapid growth of long-form video content in domains such as entertainment, education, and surveillance poses significant challenges for automated understanding systems. Unlike short clips, hour-long videos demand reasoning over tens of thousands of frames and audio streams, where events unfold across extended durations and depend on subtle cross-scene relationships. The interleaving of multiple entities, dynamic interactions, and evolving multimodal signals within such content makes coherent interpretation and entity tracking particularly difficult.
@@
Recent advances in large Vision-Language Models (VLMs) have improved tasks like captioning, question answering, and temporal grounding on short videos, yet their application to long videos remains constrained by limited context windows and computational bottlenecks. To mitigate these issues, techniques such as adaptive sampling [9, 30, 12], and token compression [24, 22, 21, 15] are proposed, enabling longer sequences under resource limits. Memory-based approaches [5, 8] further extend temporal coverage by dynamically retaining and updating salient information. While these techniques improve scalability, they often sacrifice critical details or contexts, and struggle to preserve the semantic continuity across distant segments.
To address semantic continuity, two new paradigms have emerged: retrieval-augmented generation (RAG) and video agents. RAG-based approaches segment videos into retrievable units and dynamically fetch relevant content during inference, enabling scalable and context-aware comprehension [19, 13, 10, 29, 20]. In parallel, agent frameworks treat large language models as autonomous agents capable of planning, querying, and reasoning over video content [25, 4, 31]. Despite recent progress, RAG-based and agent-based paradigms still face fundamental limitations when applied to long-form video understanding. First, retrieval often depends on local cues such as clip-level captions, producing fragmented and redundant information that undermines semantic coherence. Second, the absence of a hierarchical representation of video content prevents structured reasoning for agents. Finally, without global context, models can only resort to multiple rounds of retrieval to reconstruct continuity, introducing unnecessary complexity.
@@
In this paper, we overcome these limitations by proposing a unified, agentic framework for long video understanding that transitions from fragmented processing to coherent comprehension (see Fig. 1). Our approach is based on two core innovations: audiovisual entity cohesion and hierarchical indexing. First, we introduce a mechanism for entity-level cohesion across visual and auditory modalities, enabling consistent tracking and integration of entities throughout the video. This mitigates semantic fragmentation and preserves narrative continuity. Second, we construct a hierarchical information database that organizes video content at multiple levels—global summary, scene, segment, and entity—allowing for multi-granularity retrieval. Built atop this index, our agentic search mechanism dynamically navigates the hierarchy, enabling goal-driven reasoning. By integrating these components, our framework achieves a holistic understanding of long videos, effectively bridging fragmented information, preserving semantic consistency, and enabling scalable, structured reasoning.
Our contributions are summarized as follows:
@@
  • Audiovisual entity cohesion: we propose a cross-modal entity integration mechanism that maintains semantic consistency across time and modality, improving entity tracking and narrative coherence.
@@
  • Hierarchical indexing with agentic search: we design a multi-level indexing database and apply an agentic search strategy that dynamically queries and reasons over this hierarchy, enabling effective information access.
@@
  • We validate our framework on long video understanding benchmarks, showing superior performance compared to existing baselines. Notably, our method achieves an accuracy of 81.0% on the LVBench benchmark.
@@ 2 Related Work
RAG has emerged as a promising paradigm for scaling long video understanding by segmenting videos into retrievable units and dynamically fetching relevant context during inference. Recent works have advanced this paradigm in several directions. VideoRAG [10] retrieves aligned video-text chunks to improve factual grounding. VideoRAG [19] explores graph-based RAG over video captions to enhance entity-level semantics. AdaVideoRAG [29] introduces omni-contextual adaptive retrieval, enabling flexible selection of relevant segments through query intent classification. Vgent [20] further develops video graphs that capture temporal dependencies across clips and support structured reasoning over retrieved content. Although RAG-based methods improve retrieval precision over LVLMs, their reliance on fragmented segments and lack of global context limit temporal coherence, thus making complex reasoning tasks challenging.
@@
Unlike static RAG-based retrieval, recent work leverages LLMs as autonomous agents for iterative planning, retrieval, and reasoning over long videos, enabling dynamic interaction through tool use and structured search. Representative works include VideoAgent [25], which treats LLMs as agents that iteratively retrieve and interpret video segments; VideoTree [26], which introduces a tree-based representation for adaptive exploration; LVAgent [4], which coordinates multiple agents in multi-round collaboration; and Deep Video Discovery (DVD) [31], which supports tool-augmented search over global and local video content. Additionally, DrVideo [14] reformulates long videos into document-like structures for iterative retrieval-augmented inference. While these frameworks improve interactive reasoning and scalability, their underlying databases remain simple (e.g., frames, clip captions, visual entities) and often require heavily iterative processes to find the answer, leading to high computational cost.
@@
Long video understanding often requires modeling local-global context and hierarchical relationships across scenes, segments, and entities. Prior works address this via local-global aggregation [27, 18], tree-based structure [26], and graph-based retrieval [29, 19], while [31] tracks subjects in a global registry. Although these methods improve semantic coherence, some models incur high online computation and lack a unified offline hierarchical index spanning video, scene, segment, and entity levels—a core feature of our framework for efficient and coherent long video reasoning.
@@ 3.1 Overview
Recent RAG-based and agent-based frameworks for long-video understanding typically construct their databases using clip-level captions [31] and global entity sets that link disjoint clips [19, 29, 20]. However, this combination of local-clip and global-entity structure does not adequately reflect the inherent hierarchical nature of video content, where scenes, events and entities appear and evolve in alternating patterns. For global or scene-level queries, e.g. “what is the video about?” or “What is the name of the song sung by the third contestant?”, as well as locally ambiguous queries like “what does the protagonist do at 12:00-12:30?", models require long-range contextual information. Relying solely on online retrieval of local fragments can result in incomplete answers or overwhelm the model with redundant and incoherent information. To address this, we propose an offline hierarchical video indexing approach that parses the video into structured representation at the global, scene, segment and entity levels prior to querying. Furthermore, most existing methods focus primarily on the visual modality and may use subtitles as supplementary text, while overlooking the role of speaker identity. Unlike visual appearance or behavior, which may vary across clips, speaker voice remains relatively consistent, offering a robust cue for maintaining long-range identity coherence.
To overcome these limitations, we introduce a framework for hierarchical video indexing and agentic retrieval, which integrates audiovisual entity cohesion and multi-level temporal abstraction, enabling coherent and efficient long video understanding. We aim to perform an offline parsing to construct a structured database with four-level hierarchy:
@@
where $\tilde{\mathcal{C}}$, $\tilde{\mathcal{E}}$, $\tilde{\mathcal{S}}$, and $\tilde{\mathcal{G}}$ denote the audiovisual segment-level information, canonical audiovisual entities, scene summaries, and global summary, respectively, as shown in Figure 1. Specifically, speaker diarization is leveraged to maintain consistent speaker identities in audio modality across the entire video. Such speaker identities are then fused with visual modality in a local-to-global manner, forming coherent audiovisual entities that align speech and visual appearance across the entire video. Meanwhile, a LLM-based temporal summarization pipeline merges consecutive semantically related segments into scene-level and global abstractions, providing higher-level contextual anchors for global reasoning.
During inference, we employ an agentic reasoning and retrieval approach that performs query-dependent adaptive search over the video indexing $\mathcal{D}$ through a reasoning-driven loop: $\textit{Think}\rightarrow\textit{Act}\rightarrow\textit{Observe}$ (see Figure 1), following the paradigm introduced in [31]. We introduce a suite of multi-level retrieval tools for querying $\mathcal{D}$. Depending on the granularity and complexity of each query, the planner dynamically selects appropriate tools, ranging from global scene browsing, entity querying to local fine-grained search and inspection, to gather the necessary information for accurate response generation.
Figure 2: Overview of our framework. Left: hierarchical database construction. Right: agentic reasoning. The reasoning LLM calls tools iteratively to collect information and answer the question.
@@ 3.2.1 Audio Information Extraction
Audio serves as a crucial carrier of semantic information, conveying narrative flow, speaker identity, and contextual cues that often extend beyond the visual modality. While most existing works merely utilize transcripts or subtitles as textual supplements, they often overlook the long-range consistency provided by speaker identity. Unlike visual appearance or actions, which may vary significally across scenes, speaker voice remains relatively stable, offering a reliable signal for maintaining temporal coherence, especially in dialog-driven content such as documentaries, TV shows, and vlogs.
To leverage this, we employ automatic speech recognition (ASR) and speaker diarization using WhisperX [3], which jointly produce accurate transcripts with timestamps and consistent speaker labels. These annotations enable us to understand not only “what was said and when”, but also “who said it”, a key factor for maintaining entity-level coherence across time.
@@ 3.2.2 Segment Information Extraction
We uniformly divide the entire video into fixed-length temporal segments, which serve as the basic unit of local information. For each video segment $i$, we first extract the corresponding audio content, including speaker labels $P_{i}$ and timestamped transcripts $T_{i}$. These audio features, along with the video frames of the segment, are fed into a VLM to generate a segment-level caption $V_{i}$ and fused speaker information $P_{i}^{\prime}$ that associates visual cues with speaker labels (e.g., appearance, actions, identity for a speaker label). We then construct the textual representation of the segment $C_{i}^{t}$ by concatenating the fused speaker information $P_{i}^{\prime}$, transcript $T_{i}$, and visual caption $V_{i}$:
@@
This representation serves as the foundation for subsequent entity-level cohesion and cross-temporal reasoning.
As captioning alone may omit critical visual details, we further incorporate a visual embedding $C_{i}^{v}$ using the universal multimodal retrieval method UNITE [11] to enrich the representation. The final unified segment-level representation is defined as
@@
which grounds each utterance and provides localized visual context for each speaker, composing the segment database $\tilde{\mathcal{C}}$. In our implementation, each segment spans 30 seconds, and we sample 20 frames per segment for caption generation.
@@ 3.2.3 Audiovisual Entity Extraction
For each segment $i$, we employ a LLM to extract key entities from its textural representation $C_{i}^{t}$. The LLM identifies and categorizes salient subjects like characters, locations, and events, producing a set of entities with summarized descriptions:
@@
where each $e_{k}^{i}$ represents a semantic entity, including its name and description derived from the audiovisual context of segment $i$.
While prior approaches often merge entities using simple name matching or cosine similarity thresholds [19, 20], such heuristics are insufficient for long videos with complex variations. For example, a character appearing in different scenes with varying outfits (e.g. red dress vs. yellow dress) or emotional states (e.g. happy old man vs. angry male teacher) may be incorrectly split into multiple entities, while two visually similar but distinct characters may be mistakenly merged.
To address this, we perform entity consolidation in two stages, embedding-based clustering followed by LLM-based canonicalization. First, all entity descriptions $\{e_{k}^{i}\}$ are embedded using a text encoder $f_{\text{text}}(\cdot)$ to obtain dense representations: $z_{k}^{i}=f_{\text{text}}(e_{k}^{i}).$ We then apply clustering to group semantically similar entities into preliminary clusters, capturing potential cross-segment correspondences. Second, each cluster is re-evaluated by an LLM, which reads all member entities and either generates a unified canonical summary or separates them into distinct subgroups if semantic conflicts are detected. This LLM-in-the-loop verification prevents both over-merging and over-splitting, resulting in a refined set of canonical entities:
$\tilde{\mathcal{E}}_{g}=\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{J}\},$ (5)
where each $\tilde{e}_{j}$ aggregates temporally distributed mentions of the same underlying identity or concept.
To further enhance coherence, we emphasize speaker label cues into the consolidation process. If multiple segments share the same speaker label, their corresponding character entities about characters are prioritized for merging, even when visual or textual features differ. This enables robust audiovisual entity cohesion, ensuring consistent identity tracking across the video. For instance, a host narrating in both indoor and outdoor scenes will be recognized as the same entity based on voice consistency, rather than being fragmented due to appearance changes.
After consolidation, each canonical entity $\tilde{e}_{j}$ is associated with a global description and a set of linked segments $Q_{j}$. During retrieval, incorporating all linked segments of top-$K_{1}$ entities can be computationally expensive. Moreover, these segments often contain extraneous query-irrelevant details, degrading cosine similarity matching between the query and segment embeddings, while LLM-base reranking incurs high computational cost. To address this, we further introduce an entity-centric re-captioning process on original segment textual representation during offline database construction. For each linked segment $i$ and entity $\tilde{e}_{j}$, we generate a focused description $\tilde{C}^{t}_{i,j}$ using an LLM. This description centers on the entity’s appearance, actions, and events within the segment, while excluding irrelevant context. The final entity database consists of $\tilde{\mathcal{E}}_{g}$ and this fine-grained entity-segment links $\tilde{\mathcal{E}}_{e}=\{\tilde{C}^{t}_{i,j}|j=1,2,\ldots,J,i\in Q_{j}\}$, i.e. $\tilde{\mathcal{E}}=\{\tilde{\mathcal{E}}_{g}\,;\,\tilde{\mathcal{E}}_{e}\}$. During retrieval, after initial entity matching in embedding space of $\tilde{\mathcal{E}}_{g}$, we re-rank candidate segments by computing similarity between the query and $\tilde{C}^{t}_{i,j}$, selecting the top-$K_{2}$ segments for precise entity grounding. This approach balances retrieval accuracy and computational efficiency, avoiding redundant noise and excessive LLM overhead.

### v2

@@ 1 Introduction
The rapid growth of long-form video content in domains such as entertainment, education, and surveillance poses significant challenges for automated understanding systems. Unlike short clips, hour-long videos demand reasoning over tens of thousands of frames and audio streams, where events unfold across extended durations and depend on subtle cross-scene relationships. The interleaving of multiple entities, dynamic interactions, and evolving multimodal signals makes coherent interpretation and entity tracking particularly difficult.
@@
Recent advances in large Vision-Language Models (VLMs) have led to substantial progress on short-video tasks such as captioning, question answering, and temporal grounding, yet their application to long videos remains constrained by limited context windows and computational bottlenecks. To mitigate these issues, techniques such as adaptive sampling [9, 30, 12], and token compression [24, 22, 21, 15] are proposed, enabling longer sequences under resource limits. Memory-based approaches [5, 8] further extend temporal coverage by dynamically retaining and updating salient information. While these techniques improve scalability, they often sacrifice critical details or contexts, and struggle to preserve the semantic continuity across distant segments.
To address these issues, recent methods increasingly adopt retrieval-augmented generation (RAG) [19, 13, 10, 29, 20] to dynamically fetch relevant video segments, alongside agentic frameworks [25, 4, 31] that autonomously plan and reason over the video content. However, both paradigms exhibit fundamental limitations in long-video scenarios. First, retrieval is typically driven by isolated signals (e.g., clip-level captions), which yields fragmented or redundant evidence and severely weakens global narrative coherence. Second, the absence of a hierarchical video representation deprives agents of the structural context needed for multi-level reasoning. Consequently, models can only resort to inefficient, multi-round retrievals to recover cross-segment continuity, introducing unnecessary complexity.
@@
In this paper, we propose a unified agentic framework that shifts long-video understanding from fragmented retrieval to coherent, structured comprehension (see Fig. 1). Our approach is based on two core innovations: audiovisual entity cohesion and hierarchical indexing. First, we introduce audiovisual entity cohesion, which leverages complementary audio and visual evidence to consolidate fragmented observations into consistent entities and enrich them with multimodal attributes. Particularly, we explicitly exploit speaker identity cues as a powerful yet often over-looked signal for entity consolidation, serving as an effective “glue” for maintaining long-range entity coherence. The resulting entity representations provide reliable building blocks for higher-level scene interpretation as well. Second, we construct a hierarchical database that organizes video content at multiple granularities, i.e. global summary, scene, segment, and entity, enabling flexible and multi-level retrieval. Built atop this hierarchy, our agentic search mechanism supports goal-driven reasoning across granularities. Ultimately, by integrating these components, our framework achieves scalable and holistic long-video understanding. Our contributions are summarized as follows:
@@
  • Audiovisual entity cohesion: we propose a cross-modal entity consolidation mechanism that maintains semantic consistency across time and modalities, effectively improving entity continuity and narrative coherence.
@@
  • Hierarchical indexing with agentic search: we design a hierarchical indexing database with an agentic search strategy that dynamically navigates and reasons over this hierarchy, enabling efficient and structured information access.
@@
  • State-of-the-art performance: we evaluate our framework on several long video understanding benchmarks, achieving superior performance over existing baselines, with an overall accuracy of 84.1% on LVBench.
@@ 2 Related Work
RAG has emerged as a promising paradigm for scaling long video understanding by segmenting videos into retrievable units and dynamically fetching relevant context during inference. Recent works have advanced this paradigm in several directions, including aligned video-text chunking [10], graph-based entity semantics [19], omni-contextual adaptive selection [29], and temporal dependency modeling [20]. However, their reliance on fragmented segments and lack of global context limit temporal coherence, thus making complex reasoning tasks challenging.
@@
Unlike static RAG-based retrieval, recent work leverages LLMs as autonomous agents for iterative planning, retrieval, and reasoning over long videos, enabling dynamic interaction through tool use and structured search. Representative works include VideoAgent [25], which treats LLMs as agents that iteratively retrieve and interpret video segments; VideoTree [26], which introduces a tree-based representation for adaptive exploration; LVAgent [4], which coordinates multiple agents in multi-round collaboration; and Deep Video Discovery (DVD) [31], which supports tool-augmented search over global and local video content. Additionally, DrVideo [14] reformulates long videos into document-like structures for iterative retrieval-augmented inference. While these frameworks improve interactive reasoning and scalability, their underlying databases remain simple (e.g., frames, clip captions, visual entities) and often require heavily iterative processes to find the answer, leading to high computational costs.
@@
Long video understanding often requires modeling local-global context and hierarchical relationships across scenes, segments, and entities. Prior works address this via local-global aggregation [27, 18], tree-based structure [26], and graph-based retrieval [29, 19], while [31] tracks subjects in a global registry. Although these methods improve semantic coherence, some models incur high online computation and lack a unified offline hierarchical index spanning video, scene, segment, and entity levels, which is a core feature of our framework for efficient and coherent long video reasoning.
Figure 2: Overview of our framework. Left: hierarchical database. Right: agentic reasoning. The reasoning LLM calls tools iteratively to collect information and answer the question.
@@ 3.1 Overview
Recent RAG and agentic frameworks typically construct databases from clip-level captions [31] and global entity sets that link disjoint clips [19, 29, 20]. However, this combination of local-clip and global-entity does not capture the inherent hierarchical nature of video content, where entities, events, and scenes evolve at different temporal scales and interact over long horizons. Consequently, both global/scene-level queries (e.g. “what is the video about?” or “What is the name of the song sung by the third contestant?”) and locally ambiguous queries (e.g., “what does the protagonist do at 12:00-12:30?”) require long-range contextual reasoning. Relying solely on online retrieval of local fragments can yield incomplete evidence or overwhelm the model with redundant and incoherent information.
To address these issues, we propose a framework for semantic-coherent hierarchical video indexing and agentic retrieval that performs offline parsing to construct a structured database and enables query-dependent navigation over multiple granularities. A key ingredient is audiovisual entity cohesion: rather than treating audio as subtitle text only, we leverage speaker identity as a complementary and often more stable cue for entity continuity. Speaker identity can remain informative even when visual evidence becomes unreliable due to occlusion, changes in viewpoint or lighting, motion blur, crowded scenes, shot transitions, or off-screen speakers. This cross-modal cue helps consolidate fragmented entity observations and provides more reliable building blocks for higher-level scene interpretation.
Concretely, we construct a four-level hierarchical database
@@
where $\tilde{\mathcal{C}}$, $\tilde{\mathcal{E}}$, $\tilde{\mathcal{S}}$, and $\tilde{\mathcal{G}}$ denote segment-level audiovisual information, canonical audiovisual entities, scene summaries, and a global summary, respectively (Fig. 1). We leverage speaker diarization to maintain consistent speaker identities across the entire video, which are used together with visual information to support robust entity consolidation. Meanwhile, a LLM-based temporal abstraction pipeline merges semantically related segments into scene-level and global summaries, providing higher-level anchors for long-range reasoning.
During inference, our agent performs query-dependent adaptive search over $\mathcal{D}$ via a think-act-observe loop (Fig. 1), following [31]. A suite of multi-granularity retrieval tools allows the agent to dynamically navigate at multiple levels, enabling efficient evidence collection and accurate answer generation for diverse query types.
@@ 3.2.1 Audio Information Extraction
Audio provides complementary cues for long-video understanding, including time-aligned transcripts and speaker identity. Beyond using transcripts/subtitles as auxiliary text, we incorporate speaker identity as a long-range consistency signal that remains informative even when visual cues degrade (e.g., occlusion, shot transitions, view changes, or off-screen speakers). This is particularly valuable for dialog-driven content such as documentaries, TV shows, and vlogs. We employ automatic speech recognition (ASR) and speaker diarization using WhisperX [3], which jointly produce accurate transcripts with timestamps and consistent speaker labels. These annotations enable us to understand not only “what was said and when” but also “who said it,” a key factor for maintaining entity-level coherence across time.
@@ 3.2.2 Segment Information Extraction
We uniformly divide the video into fixed-length temporal segments as the basic unit of local evidence. For each segment $i$, we extract the audio annotations, including speaker labels $P_{i}$ and timestamped transcripts $T_{i}$. These audio annotations, together with sampled video frames, are provided to a VLM to produce a segment-level caption $V_{i}$ and a speaker-aware description $P_{i}^{\prime}$, which associates each speaker label with salient visual cues in the segment (e.g., appearance, actions, role cues when available). We then construct the segment textual representation $C_{i}^{t}$ by
@@
This representation grounds who said what and when in a localized visual context and serves as the basis for subsequent entity consolidation and cross-temporal reasoning.
Since captioning may miss fine-grained visual details, we further augment each segment with a visual embedding $C_{i}^{v}$ using the multimodal retrieval model UNITE [11]. The final segment representation is
@@
which composes the segment database $\tilde{\mathcal{C}}$. In our implementation, each segment spans 30 seconds and we sample 20 frames per segment for caption generation.
Figure 3: Speaker identity for entity consolidation. Segments with long temporal gaps can still be linked to the same speaker entity through a shared speaker identity.
@@ 3.2.3 Audiovisual Entity Extraction
For each segment $i$, we employ an LLM to extract salient entities from the textual representation $C_{i}^{t}$, covering characters, locations, and events. This yields
@@
where each $e_{k}^{i}$ includes an entity name and a concise description derived from the segment’s audiovisual context.
Consolidating these entities across long videos is non-trivial. Simple heuristics such as name matching or embedding-similarity thresholding [19, 20] can be brittle: the same character may be split across segments due to appearance or viewpoint changes, whereas distinct but visually similar characters may be erroneously merged. To address this, we perform entity consolidation in two stages, embedding-based clustering followed by LLM-based canonicalization. First, we embed each entity description using a text encoder $f_{\text{text}}(\cdot)$, $z_{k}^{i}=f_{\text{text}}(e_{k}^{i})$, and cluster the embeddings to form candidate groups of cross-segment correspondences. Second, an LLM revisits each cluster and either produces a canonical entity summary or splits the cluster into subgroups when semantic conflicts are detected. This verification step mitigates both over-merging and over-splitting, yielding a refined set of canonical entities
$\tilde{\mathcal{E}}_{g}=\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{J}\}.$ (5)
To further strengthen long-range coherence, we incorporate speaker identity cues into consolidation. When multiple segments share the same diarized speaker label, we treat this as a strong consistency signal and prioritize merging the corresponding character-related entity mentions, even when visual or textual descriptions vary due to shot changes, occlusion, or other degradations. This yields robust audiovisual entity cohesion and improves identity continuity across the video. Figure 3 illustrates two examples in which characters undergoing dramatic appearance changes are correctly consolidated using speaker identity cues. Such consolidation is crucial for questions like “How does the emotion change on Sarah’s face when interviewed?”, where speaker identity links disjoint fragments to the identity “Sarah”, capturing a long-range continuity that isolated clip captions fail to provide.
After consolidation, each canonical entity $\tilde{e}_{j}$ is associated with a global description and a set of linked segments $Q_{j}$. Directly incorporating all segments linked to the top-$K_{1}$ entities can be expensive and may introduce query-irrelevant noise. We therefore perform entity-centric re-captioning during offline construction, producing a focused description $\tilde{C}^{t}_{i,j}$ centering on the entity $\tilde{e}_{j}$ for each linked segment $i\in Q_{j}$. The final entity database contains both canonical entities and fine-grained entity-segment descriptions, $\tilde{\mathcal{E}}=\{\tilde{\mathcal{E}}_{g}\,;\,\tilde{\mathcal{E}}_{e}\}$, where $\tilde{\mathcal{E}}_{e}=\{\tilde{C}^{t}_{i,j}|j=1,2,\ldots,J,i\in Q_{j}\}$. More details on this process are provided in Sec. A of the supplementary material.